# Identifying Spurious Correlations using Counterfactual Alignment

### v2

@@ Abstract
Models driven by spurious correlations often yield poor generalization performance. We propose the counterfactual (CF) alignment method to detect and quantify spurious correlations of black box classifiers. Our methodology is based on counterfactual images generated with respect to one classifier being input into other classifiers to see if they also induce changes in the outputs of these classifiers. The relationship between these responses can be quantified and used to identify specific instances where a spurious correlation existsThThis is validated by observing intuitive trends in a face-attribute face-attribute and waterbird classifiers, as well as by fabricating spurious correlations and detecting their presence, both visually and quantitatively. Furthermore, utilizing the CF alignment method, we demonstrate that we can evaluate robust optimization methods (GroupDRO, JTT, and FLAC) by detecting a reduction in spurious correlations.
00footnotetext: AWork not related to position at Amazon.
@@ 1 Introduction
Challenges related to neural network generalization and fairness often arise due to covariate shift (Moreno-Torres et al., 2012) and shortcut learning (Ross et al., 2017; Geirhos et al., 2020). Shortcut learning can lead to models making decisions based on factors not aligned with expectations of the model creator. These powerful models (vision models in this work) leverage a wide array of features and relationships, which may inadvertently incorporate unwanted spurious correlations. These spurious relationships may stem from sample bias (e.g., predicting cows when cows are observed on grass but not on a beach) or may be inherent to the class definition (e.g., predicting cows when an animal has four legs) (Beery et al., 2018).
In this work our objective is to understand black-box classifiers, without access to their training data, as this is a common use case encountered by practitioners. In this analysis, we utilize counterfactual (CF) images which are synthetic images simulating a change in the class label of an image (Pearl, 2009). These synthetic images have features modified such that the prediction of the classifier changes. We can then view the synthetic images to understand the reasons that a prediction was made.
Specifically, we are interested in CF images that are directly generated using the gradients of a classifier (Cohen et al., 2021; Joshi et al., 2018; Balasubramanian et al., 2020). Generating CF images with respect to a classifier is rooted in similar logic to that of crafting adversarial examples. However, the key distinction lies in the constraint that CF images remain within the data manifold of plausible images. The latent space of an autoencoder provides such data manifold. This approach enables us to study the specific features used by a given classifier on a particular input. By keeping the classifier and autoencoder independent, different classifiers can be analysed using a singular fixed autoencoder as a reference point.
@@
The approach we take is to generate CF images for the positive predictions of a classifier over an entire dataset and study the resulting CF images using a collection of different classifiers, which we refer to as downstream classifiers, to observe what outputs are impacted. If the CF images generated for one classifier also change the predictions of other classifiers, we can conclude that there is shared feature usage. This shared feature usage can prompt us to investigate unexpected relationships. The relative change metric (Eq. 3) can be used to quantify the impact of counterfactual perturbations and identify spurious correlations that may occur for a specific example or to identify trends across an entire dataset. We study face-attribute classification, recognizing its advantage in allowing visual verification by readers.
@@
  • We propose CF alignment to reason about the feature relationships between classifiers, quantified using the relative change metric. This approach allows for both aggregate quantification and targeted querying of models to locate specific examples where the spurious correlations are used to make predictions.
@@ 3 CF Alignment Methodology
In this work, we introduce the CF alignment approach, outlined in Figure 1. The methodology involves the generation of counterfactual samples using a technique known as Latent Shift (Cohen et al., 2021). The implementation of Latent Shift requires the integration of an encoder/decoder model, denoted as $D(E(x))$, where $E$ represents the encoder and $D$ is the decoder. Additionally, a classifier $f$ is incorporated, responsible for predicting the target variable $y$, expressed as $y=f(x)$. It is important to note that both the autoencoder and the classifier are trained independently, with the only specified requirements being differentiability and operation on the same data domain.
@@ 3.1 Relative Change
$R(f_{1},f_{b},z_{0})=\frac{f_{1}(D(z_{\lambda}))-f_{1}(D(z_{0}))}{f_{b}(D(z_{\lambda}))-f_{b}(D(z_{0}))}$ (3)
@@
In our experiments, we opted for relative change over traditional correlation measures. This decision was motivated by the observation that correlation metrics occasionally yielded false positives when only a slight change in the prediction of $f_{1}$ occurred compared to the base classifier $f_{b}$. Relative change provides a more nuanced understanding by considering not just the direction but also the magnitude of the change, offering a more robust assessment of the impact of counterfactual perturbations on downstream classifiers.
The relative change metric is a key contribution that allows us to make sense of the CF alignment results at scale over large datasets.
@@ 4 Experiments
The pre-trained face classifiers used in this work are sourced from (Vandenhende et al., 2020). They were trained on the CelebA dataset (Liu et al., 2015) to predict 40 different facial attributes on images of dimensions 224$\times$224. We chose this model because it has good performance, is publicly available, and was implemented in PyTorch.
We leverage the VQ-GAN autoencoder from (Esser et al., 2021) trained on the FacesHQ dataset, which combines the CelebA HQ dataset (Karras et al., 2018) and the Flickr-Faces-HQ (FFHQ) dataset (Karras et al., 2021). The resolution of this model is 256x256. In order to make this work for existing classifiers, each classifier normalizes the input image dynamically to match its training domain.
The Captum library (Kokhlikyan et al., 2020) is utilized for baseline attribution methods. PyTorch (Paszke et al., 2019) is used for efficient tensor computation and automatic differentiation. The source code and model weights for all experiments will be released publicly online [^1] https://github.com/ieee8023/latentshift. The CF alignment algorithm requires a few seconds to run for each image on a NVIDIA V100 16GB GPU. The CF generation step is the most expensive and is variable due to the automated search for the optimal lambda. The resulting CF images are then processed by each classifier which scales by the number of classifiers studied.
@@ 4.1 Aggregate statistics over a dataset
Viewing aggregate CF alignment statistics over a dataset can be useful when investigating a model for bias or spurious correlations. The average relative change between pairs of classifiers (N=400 per class) is shown in Figure 2a. Figures 2b and 2c show correlations between classifier predictions and the ground truth training labels. A subset of classifiers studied are shown in this figure for clarity, with the full CF alignment matrix in Appendix Fig. 7. These classifiers were chosen because they have intuitive and unintuitive relationships with large relative change that help to illustrate the contribution of this method.
The caption of Figure 2 details observations where correlation in the training data and classifier outputs are not the same as the relationships uncovered using CF alignment. These indicate that even if correlated attributes are presented to the classifier it does not cause feature usage to be correlated. These observations also imply that even if uncorrelated attributes are presented to the classifier it may still construct spurious feature relationships.
@@ 4.2 Studying specific examples
Figure 3: CF alignment examples for pointy_nose with the highest aligned and inverse aligned classifiers. We observe an inverse alignment with big_nose and potential spurious relationships with eyebrows, eyes, hair, and smiling. The relative change is shown next to each classifier name.
Using the aggregate analysis from §4.1 as a guide, classifiers can be selected to investigate undesired relationships on specific images. Inspecting the pointy_nose classifier in Figure 3 using two images with positive predictions for pointy_nose we observe similar and contrasting relationships between classifiers. A common theme is the inverse alignment with big_nose which is intuitive as it is the opposite of a pointy nose. rosy_cheeks is a common aligned classifier which does not appear to have an obvious reason and is likely a spurious correlation. An alignment with smiling is only observed in one of the examples.
@@
4.3 Comparison with saliency maps
Saliency maps for feature attribution could potentially be used to perform a similar analysis by looking at relationships between their generated heatmaps. Here, we present an example that demonstrates their limitation. The attribution methods input gradients, integrated gradients Sundararajan et al. (2017), and guided backprop Springenberg et al. (2015) are used to explain the classifier predictions of pointy_nose and big_nose.
@@ 4.4 Validation by inducing spurious correlations
(a) CF for the classifier smiling showing eyebrows are unchanged. The horizontal line indicates the prediction of arched_eyebrows is not influenced by the features used for smiling in this image.
@@
Figure 5: Example of detecting a spurious correlation in a biased classifier. The classifier is biased with arched eyebrows and this is observed in the alignment plot as well as in the counterfactual image. The relative change is now 0.97 compared to 0.01 for the unchanged smiling classifier.
@@ 4.5 Analysis of robust optimization methods
Experiments in this section are performed on the held out test datasets and selected to be balanced such that there is a balanced distribution of positive and negative examples with the sensitive attribute (e.g. in the CelebA dataset, 1/4 of the samples are labeled to have Blond Hair and be Male, 1/4 Blond Hair and be Not Male, 1/4 Not Blond Hair and be Not Male, and 1/4 Not Blond Hair and be Male).
During classification experiments, the samples evaluated are selected to be inversely correlated with the sensitive attribute to cause the most discrepancy in performance (e.g. Samples are selected that are labeled Blond Hair and Male as well as Not Blond Hair and Not Male).
@@ 4.5.1 GroupDRO on Waterbirds
Three models are evaluated for this experiment, a baseline model which was induced to rely on spurious correlations, a model trained with Group Distributionally Robust Optimization (DRO) (Sagawa et al., 2020) which aims to minimize error over groups (which are known during training, in this case the background class), a model using Just Train Twice (JTT) (Liu et al., 2021) which uses a two stage approach that boosts misclassified training examples in a second training cycle. The classifiers that these models are aligned with are trained to predict if the background of an image is land or water using the same training dataset labels from the Waterbirds dataset.
@@ 5 Limitations
There are limitations and challenges to CF generation Verma et al. (2020) where bias can also exist in the CF generation method. Furthermore, although we are optimistic that our method will generalize to additional domains, we focus our current analysis on face-attribute classification. Although we do not demonstrate the generalizability of our method to additional domains, using face-attribute classification is a deliberate choice enabling qualitative evaluation, in addition to our quantitative evaluation. Additionally, since CF generation relies on a separate autoencoder, improving the representational capability of the autoencoder may improve the fidelity of the CF generation.
@@ Appendix A Appendix
A.1 Rectifying bias by composing classifiers

### v3

@@ Abstract
Models driven by spurious correlations often yield poor generalization performance. We propose the counterfactual (CF) alignment method to detect and quantify spurious correlations of black box classifiers. Our methodology is based on counterfactual images generated with respect to one classifier being input into other classifiers to see if they also induce changes in the outputs of these classifiers. The relationship between these responses can be quantified and used to identify specific instances where a spurious correlation existsThThis is validated by observing intuitive trends in face-attribute and waterbird classifiers, as well as by fabricating spurious correlations and detecting their presence, both visually and quantitatively. Furthermore, utilizing the CF alignment method, we demonstrate that we can evaluate robust optimization methods (GroupDRO, JTT, and FLAC) by detecting a reduction in spurious correlations.
00footnotetext: Work not related to position at Amazon.
@@ 1 Introduction
Challenges related to neural network generalization and fairness often arise due to covariate shift (Moreno-Torres et al., 2012) and shortcut learning (Ross et al., 2017; Geirhos et al., 2020). Shortcut learning can lead to models making decisions based on factors not aligned with expectations of the human that created the model. These powerful models (vision models in this work) leverage a wide array of features and relationships, which may inadvertently incorporate unwanted spurious correlations. These spurious relationships may stem from sample bias (e.g., predicting cows when cows are observed on grass but not on a beach) or may be inherent to the class definition (e.g., predicting cows when an animal has four legs) (Beery et al., 2018).
In this work our objective is to understand black-box classifiers, without access to their training data, as this is a common use case encountered by practitioners. In this analysis, we utilize counterfactual (CF) images which are synthetic images simulating a change in the class label of an image (Wachter et al., 2017). These synthetic images have features modified such that the prediction of the classifier changes. We can then view the synthetic images to understand the reasons that a prediction was made.
Specifically, we are interested in CF images that are directly generated using the gradients of a classifier (Cohen et al., 2021; Joshi et al., 2018; Balasubramanian et al., 2020). Generating CF images with respect to a classifier is rooted in similar logic to that of crafting adversarial examples. However, the key distinction lies in the constraint that CF images remain within the data manifold of plausible images. The latent space of an autoencoder provides such data manifold. This approach enables us to study the specific features used by a given classifier on a particular input. By keeping the classifier and autoencoder independent, different classifiers can be analysed using a single autoencoder as a reference point.
@@
The approach we take is to generate CF images for the positive predictions of a classifier over an entire dataset and study the resulting CF images using a collection of different classifiers, which we refer to as downstream classifiers, to observe what outputs are impacted. If the CF images generated for one classifier also change the predictions of other classifiers, we can conclude that there is shared feature usage. This shared feature usage can prompt us to investigate if this relationship is unexpected. Such as if a classifier that predicts big noses should share features with a classifier that predicts arched eyebrows.
The relative change metric (Eq. 3) can be used to quantify the impact of counterfactual perturbations and identify spurious correlations that may occur for a specific example or to identify trends across an entire dataset. We study face-attribute classification, recognizing its advantage in allowing visual verification by readers.
@@
  • We propose CF alignment to reason about the feature usage relationships between classifiers, quantified using the relative change metric. This approach allows for both aggregate quantification and targeted querying of models to locate specific examples where the spurious correlations are used to make predictions.
@@ 3 CF Alignment Methodology
In this work, we introduce the CF alignment approach, outlined in Figure 1. The CF alignment approach tests if a base classifier $f_{b}$ utilizes features that are also used by a downstream classifier (e.g. $f_{1}$). If they do, the generated CF would have features modulated which would cause the other classifier (e.g. $f_{1}$) to have a different output. This alignment can then be measured quantitatively by comparing how aligned the changes in prediction are. Here we consider a classifier as predicting a single scalar value $f(x)\in[0,1]$. A multiclass classifier can also work if we consider one class vs all other classes to create a single output value.
We say that two classifiers are aligned if both their predictions change negatively with a counterfactual generated for one of them, implying the features removed were used for both predictions. We say they are inverse aligned when one classifier’s prediction is reduced and the other’s prediction increased. We discuss how to quantify this in §3.1 but will continue on how we generate CFs.
To generate CF samples a technique known as Latent Shift (Cohen et al., 2021) is used. The implementation of Latent Shift requires the integration of an encoder/decoder model, denoted as $D(E(x))$, where $E$ represents the encoder and $D$ is the decoder. Additionally, a classifier $f$ is incorporated, responsible for predicting the target variable $y$, expressed as $y=f(x)$. It is important to note that both the autoencoder and the classifier are trained independently, with the only specified requirements being differentiability and operation on the same data domain.
@@ 3.1 Relative Change
$\text{RelativeChange}(f_{1},f_{b},z_{0})=\frac{f_{1}(D(z_{\lambda}))-f_{1}(D(z_{0}))}{f_{b}(D(z_{\lambda}))-f_{b}(D(z_{0}))}$ (3)
@@
In our experiments, we opted for relative change over traditional correlation measures. This decision was motivated by the observation that correlation metrics occasionally yielded false positives when only a slight change in the prediction of $f_{1}$ occurred compared to the base classifier $f_{b}$. Relative change provides a more nuanced understanding by considering not just the direction but also the magnitude of the change, offering a more robust assessment of the impact of counterfactual perturbations on downstream classifiers. The relative change metric value can be interpreted as follows, a value of 1 indicates a very high alignment, 0 indicates no alignment/independence, and -1 indicates a strong inverse alignment. The relative change metric is a key contribution that allows us to make sense of the CF alignment results at scale over large datasets.
@@ 4 Experiments
The pre-trained face classifiers used in this work are sourced from (Vandenhende et al., 2020). They were trained on the CelebA dataset (Liu et al., 2015) to predict 40 different facial attributes on images of dimensions 224$\times$224. These classifiers all have the same architecture, the only difference between them is the last classifier layer weights. We chose this model because it has good performance, is publicly available, and was implemented in PyTorch.
We leverage the VQ-GAN autoencoder from (Esser et al., 2021) trained on the FacesHQ dataset, which combines the CelebA HQ dataset (Karras et al., 2018) and the Flickr-Faces-HQ (FFHQ) dataset (Karras et al., 2019). The resolution of this model is 256x256. In order to make this work for existing classifiers, each classifier normalizes the input image dynamically to match its training domain.
[…]
@@ 4.1 Aggregate statistics over a dataset
Viewing aggregate CF alignment statistics over a dataset can be useful when investigating a model for bias or spurious correlations. To achieve this we compute the average relative change between pairs of classifiers (N=400 images per class) is shown in Figure 2a. Figures 2b and 2c show correlations between classifier predictions and the ground truth training labels. A subset of classifiers studied are shown in this figure for clarity, with the full CF alignment matrix in Appendix Fig. 6. These classifiers were chosen because they have intuitive and unintuitive relationships with large relative change that help to illustrate the contribution of this method. The positive relationships between smiling and mouth_slightly_open and high_cheekbones seems intuitive while the positive relationship between pointy_nose and arched_eyebrows seems unintuitive and likely due to a spurious correlations. This relationship is explored on a single example in §4.2.
The caption of Figure 2 details observations where correlation in the training data and classifier outputs are not the same as the relationships uncovered using CF alignment. These observations indicate that even if correlated attributes are presented to the classifier it does not cause feature usage to be correlated. These observations also imply that even if uncorrelated attributes are presented to the classifier it may still construct spurious feature relationships.
@@ 4.2 Studying specific examples
Figure 3: CF alignment examples for the $f_{b}$=pointy_nose with the highest aligned and inverse aligned classifiers. We observe an inverse alignment with big_nose and potential spurious relationships with eyebrows, eyes, hair, and smiling. The relative change is shown next to each classifier name.
Using the aggregate analysis from §4.1 as a guide, classifiers can be selected to investigate undesired relationships on specific images. Inspecting the pointy_nose classifier in Figure 3 using two images with positive predictions for pointy_nose we observe similar and contrasting relationships between classifiers. A common theme is the inverse alignment with big_nose which is intuitive as it is the opposite of a pointy nose. rosy_cheeks is a common aligned classifier which does not appear to have an intuitive reason (to the authors) and is likely a spurious correlation. An alignment with smiling is only observed in one of the examples.
@@ 4.3 Validation by inducing spurious correlations
(a) CF for the base classifier smiling showing eyebrows are unchanged. The horizontal line indicates the prediction of arched_eyebrows is not influenced by the features used for smiling in this image.
@@
Figure 4: Example of detecting a spurious correlation in a biased base classifier. The classifier is biased with arched eyebrows and this is observed in the alignment plot as well as in the counterfactual image. The relative change is now 0.97 compared to 0.01 for the unchanged smiling classifier.
@@ 4.4 Analysis of robust optimization methods
Experiments in this section are performed on the held out test datasets and selected to be balanced such that there is a balanced distribution of positive and negative examples with the sensitive attribute (e.g. in the CelebA dataset, 1/4 of the samples are labeled to have blond_hair and be male, 1/4 blond_hair and be not male, 1/4 not blond_hair and be not male, and 1/4 not blond_hair and be male).
During classification experiments, the samples evaluated are selected to be inversely correlated with the sensitive attribute to cause the most discrepancy in performance (e.g. Samples are selected that are labeled blond_hair and male as well as not blond_hair and not male).
@@ 4.4.1 GroupDRO on Waterbirds
Three models are evaluated for this experiment, a baseline model which was induced to rely on spurious correlations, a model trained with Group Distributionally Robust Optimization (DRO) (Sagawa et al., 2020) which aims to minimize error over groups (which are known during training, in this case the background class), a model using Just Train Twice (JTT) (Liu et al., 2021) which uses a two stage approach that boosts misclassified training examples in a second training cycle. To perform CF alignment, classifiers are trained to predict land and water using the samples from the waterbirds dataset labeled as those respective classes.
@@ 5 Limitations
There are limitations and challenges to CF generation (Verma et al., 2020) where bias can also exist in the CF generation method which can limit the features able to be modulated and prevent spurious features from being detected.
Spurious alignment: This situation can occur if the downstream classifier ($f_{1}$) also has the same spurious features as the base classifier ($f_{b}$). If this is the case, then both of their output would change if the spurious features were modulated causing these classifiers to appear aligned. For example, if a base classifier predicts cow but spuriously uses the presence of a beach to produce a low prediction. A CF generated for this classifier may change the background of a farm to a beach. If we check the alignment with a classifier that predicts chicken (but this chicken classifier also reduces its prediction if the image contains a beach) then these classifiers would appear aligned and the relative change would be positive. These classifiers are in fact aligned based on feature usage but just because they just use the same spurious features. This may be similar to a false positive but it isn’t false, the classifiers are aligned.
For this reason, in this work we reference the classifier target names (e.g. mouth_slightly_open) instead of the concept that it is assumed to represent. One solution to avoid these edge cases is to utilize multiple classifiers that are trained to predict the same concept using multiple training datasets so on average they predict the correct concept.
False negative alignment: This situation can occur if the downstream classifier ($f_{1}$) predicts using features that are "opposite" of the base classifier ($f_{b}$). For example "opposite" can mean the base classifier ($f_{b}$) predicts cow using only the presence of absence of a beach while the downstream classifier ($f_{1}$) predicts beach using only the presence of absence of a cow. The CF for the base classifier changes the background from a beach to something else but this doesn’t change the prediction of the downstream classifier ($f_{1}$) so there is no alignment, the relative change would be 0. This case could be considered a false negative because we didn’t find alignment with a beach classifier but the beach classifier was flawed. To address an issue like this we could simply inspect the CFs that are being generated with domain experts as well as vary the downstream classifier’s training data.
Furthermore, although we are optimistic that our method will generalize to additional domains, we focus our current analysis on face-attribute classification and waterbirds. Although we do not demonstrate the generalizability of our method to additional domains, using face-attribute classification is a deliberate choice enabling qualitative evaluation, in addition to our quantitative evaluation. Additionally, since CF generation relies on a separate autoencoder, improving the representational capability of the autoencoder may improve the fidelity of the CF generation.
@@ Appendix A Appendix
A.1 Comparison with saliency maps
Saliency maps for feature attribution could potentially be used to perform a similar analysis by looking at relationships between their generated heatmaps. Here, we present an example that demonstrates their limitation. The attribution methods input gradients, integrated gradients (Sundararajan et al., 2017), and guided backprop (Springenberg et al., 2015) are used to explain the classifier predictions of pointy_nose and big_nose.
A.2 Comparison with other CF generation methods
Using the biased classifier from §4.3, we evaluate whether another counterfactual (CF) generation methods can capture its bias. As a baseline, we use the Wachter et al. (2017) method, which generates a new $x^{\prime}$ that minimizes the classifier’s prediction as well as the proximity to the input $x$. The objective is as follows:
$\underset{x^{\prime}}{\text{argmin}}\ \lambda\cdot d(x,x^{\prime})+(f(x^{\prime})-y^{\prime})^{2}$ (5)
The distance function $d$ is chosen to be the Frobenius norm in order to match pixels in the image. While this approach was not designed for images it is still able to change the base classifiers prediction in our experiments. The CF generated with this method is shown in Figure 8 and does not appear to generate features that capture alignment with the classifier’s bias. This method operates in a similar way to adversarial examples which introduce imperceptible changes.
Figure 8: Comparison with the Wachter counterfactual generation method. A smiling classifier with a known bias of arched_eyebrows is used to observe if this bias can be detected. The Wachter counterfactual doesn’t appear to modulate the input image in a human interpretable way, instead many small swirls appear on the face which change the prediction for smiling but not for arched_eyebrows.
A.3 Rectifying bias by composing classifiers